OpenStudy (anonymous):

Can you help it is a practices but i need help because it will show on my grade book thank you it is lesson 4 unit 1

OpenStudy (anonymous):

1. Simplify. 5–1(3–2) (1 point) A)45 B)1/45 C)15^–3 D)15^2 ---------------- 2. Simplify. mn^-4/p^0q^-2 (1 point) A)mq^2/n^4 B)mp/n^4 C)mn^4q^2p^0 D)q^2/mn^4 ---------------- 3. Write in scientific notation. 0.0042 (1 point) A)42 × 10^2 B)4.2 × 10^4 C)4.2 × 10^3 ---------------- D).42 × 10^4 4. Write in standard notation. 6.12 × 10^3 (1 point) A)6,120 ---------------- B)612 C)61,200 D)61.2 5. Simplify. Write in scientific notation. 0.5(8 × 10^5) (1 point) A)8 × 5^5 B)4 × 5^5 C)8 × 10^5 D)4 × 10^5 ---------------- 6. Simplify. Write in scientific notation. (9 × 10^3)^2 (1 point) A)8.1 × 10^5 B)81 × 10^4 C)81 × 10^3 D)8.1 × 10^6 ---------------- 7. Evaluate 1/2a^4b^2 for a = –2 and b = 4. (1 point) A)2 ---------------- B)1/2 C)3 D)4 8. Is the following statement always, never, or sometimes true? A number raised to a negative exponent is negative. (1 point) A)always B)never C)sometimes ---------------- 9. Simplify (4xy^2)^3(xy)^5 (1 point) A)64x^8y^11 B)64x^15y^30 C)12x^2y^11 D)12x^8y^11 ---------------- 10. Simplify (2t^3)^3(0.4r)^2 (1 point) A)3.2r^2 B)1.28r^2/t^-6 ---------------- C)1.28r^2/t^9 D)0.8r^2/t^-6

OpenStudy (anonymous):

can you help me

OpenStudy (anonymous):

@MiaPonce0827

OpenStudy (anonymous):

check my answers please thanks

OpenStudy (anonymous):

i believe 8 is never, because a negative exponent becomes 1/ ex: a^-b = 1/a^b

OpenStudy (anonymous):

6 is A

OpenStudy (anonymous):

how did you get a for 6 and can you help me with 7

OpenStudy (anonymous):

bases are the same so just add 3+2 (the exp) I don't get 7,9,10 D:

OpenStudy (anonymous):

too messy

OpenStudy (anonymous):

thank you for helping me @ryanvarghese12790

OpenStudy (anonymous):

1)a. 1/45 2)q^-2n^-4m 3. Write in scientific notation. 0.0042 (1 point) 4.2 × 10–3 4. Write in standard notation. 6.12 × 103 (1 point) 6,120 5. Simplify. Write in scientific notation. 0.5(8 × 105) (1 point) 4 × 10^5 6. Simplify. Write in scientific notation. (9 × 10–3)2 (1 point) 8.1 × 10–5 7. Evaluate a–4b2 for a = –2 and b = 4. (1 point) Again, I can't understand what this question is asking. 8. Is the following statement always, never, or sometimes true? A number raised to a negative exponent is negative.

OpenStudy (anonymous):

@S17.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

don't read this it only gets you a 50% at that